# Massively Parallel Proof-Number Search for Impartial Games and Beyond

### v1

@@ Appendix C Proof of Theorem 2
For the induction base, we have $d=0$, in which case our node is the root $r$. If $r$ has not yet been solved, then the MPN is always in the subtree of $r$ and the inequalities (3) are satisfied as $pt(r)=\infty>pn(r)$, $dt(r)=\infty>dn(r)$, and $\min\{pn(r)+0,dn(r)+0\}<\infty=mt(r)$. If $r$ is solved, then obviously the MPN is not in the subtree of $r$ and the threshold inequalities are not satisfied, since $\infty<\infty$ is interpreted as false.
@@ Appendix D Parameter Analysis
We begin by recalling the final scaling efficiency of PNS-PDFPN with optimally tuned parameters, as shown in Figure 7, where we also include a logarithmic scale for a more detailed evaluation. sIn this section, we analyze the optimal parameter settings of PNS-PDFPN by examining the search overhead, defined as the ratio of node expansions performed by PNS-PDFPN to those performed by a sequential DFPN. To capture the synchronization and communication overhead resulting from frequent interactions with the master, we also measure the worker utilization, defined as the average proportion of time that workers are actively assigned to a job.
@@
Also, note that beyond a certain point, adding more workers to the algorithm stops being beneficial, as the overhead on the master grows significantly. This is evident from the low utilization of the workers. In the following paragraph, we explain how to address this limitation.

### v2

@@ Appendix C Proof of Theorem 2
For the induction base, we have $d=0$, in which case our node is the root $r$. If $r$ has not yet been solved, then the MPN is always in the subtree of $r$, and the inequalities (3) are satisfied as $pt(r)=\infty>pn(r)$, $dt(r)=\infty>dn(r)$, and $\min\{pn(r)+0,dn(r)+0\}<\infty=mt(r)$. If $r$ is solved, then obviously the MPN is not in the subtree of $r$, and the threshold inequalities are not satisfied, since $\infty<\infty$ is interpreted as false.
@@ Appendix D Parameter Analysis
We begin by recalling the final scaling efficiency of PNS-PDFPN with optimally tuned parameters, as shown in Figure 7, where we also include a logarithmic scale for a more detailed evaluation. In this section, we analyze the optimal parameter settings of PNS-PDFPN by examining the search overhead, defined as the ratio of node expansions performed by PNS-PDFPN to those performed by a sequential DFPN. To capture the synchronization and communication overhead resulting from frequent interactions with the master, we also measure the worker utilization, defined as the average proportion of time that workers are actively assigned to a job.
@@
Also, note that beyond a certain point, adding more workers to the algorithm stops being beneficial, as the overhead on the master grows significantly. This is evident from the low utilization of the workers. In the following subsection, we explain how to address this limitation.